OpenStudy (anonymous):

Hey everyone, my cousins having trouble with this problem and i'm not sure how to help. 20 lb turkey, takes 5 hours to cook, how long does it take to cook per pound. It needs to be set up algebraically.

OpenStudy (anonymous):

It takes 5 hours to cook 20 lbs. of turkey, and you need to find hw long it takes to cook 1 lb. You can set up a proportion for this

OpenStudy (anonymous):

It should take 15 minutes to cook one pound. If you go with a proportion, you can do 5/20 = x/1 and cross multiply, making it 20x=5,making x equal 1/4 which in cooking time 1/4 of an hour is 15 minutes.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

\[\frac{300 \min }{20 \text{ lb}}= 15 \min /\text{lb} \]